OpenStudy (anonymous):

FAN + MEDAL!! Which of the following points lies on the graph of y = x2 - 2x + 6? (-3, 21) (-2, 0) (-1, 10)

OpenStudy (bradely):

y = x^2 - 2x + 6 (-3,21) 21=(-3)^2-2(-3)+6 21=9+6+6 21=21 So(-3,21)is lies on the graph Source:www.mathskey.com

OpenStudy (texaschic101):

yep...thats how you would do it. sub in your points and see which ones come out equal